OpenStudy (anonymous):

Given tanA=5 and sinB=2/3 I need to find: sin(A-B) cos(A-B) tan(A-B)

OpenStudy (jdoe0001):

so.... what's stopping from getting sin(A-B)?

OpenStudy (anonymous):

Because I dont know how to find Acos and Bcos

OpenStudy (jdoe0001):

ahhh, well... ok... let's us notice something

OpenStudy (anonymous):

Ya do you know how to find that with the given information?

OpenStudy (jdoe0001):

\(\bf tan(A) = 5 \\ \quad \\ \quad \\ tan(\theta) = \cfrac{b}{a}\quad thus\\ \quad \\ tan(A) = 5\implies tan(A) = \cfrac{5}{1}\implies \cfrac{b}{a}\\ \quad \\ c^2 = a^2+b^2\implies c = \sqrt{a^2+b^2}\\ \quad \\ \textit{keep in mind that }\quad cos(A) = \cfrac{a}{c}\)

OpenStudy (anonymous):

hmm... Im still not quite understanding.

OpenStudy (jdoe0001):

the idea being that tan(A) = 5, or 5/1, keeping in mind the tangent identity of \(\bf tan(\theta) = \cfrac{\textit{opposite side}}{\textit{adjacent side}}\implies \cfrac{b}{a}\)

OpenStudy (jdoe0001):



OpenStudy (anonymous):

Okay I got that.. The what do I do?

OpenStudy (jdoe0001):

you find the "c" side of angle A, once you find "c" you find the cosine, because cos(A) = a/c

OpenStudy (anonymous):

Okay so cos(A)=1/\[\sqrt{26}\]?

OpenStudy (anonymous):

\[1/\sqrt{26}\]

OpenStudy (jdoe0001):

\(\bf cos(A) = \cfrac{1}{\sqrt{26}}\) then

OpenStudy (anonymous):

Okay so my last question is how would I find tanB?

OpenStudy (jdoe0001):

the 2nd one is a bit ambiguous, because the angle could be on either the 1st or 2nd Quadrants, but anyhow \(\bf sin(B) = \cfrac{2}{3}\implies \cfrac{b}{c}\\ \quad \\ c^2 = a^2 + b^2 \implies \sqrt{c^2-b^2}=a\\ \quad \\ \textit{keep in mind that }\quad cos(B) = \cfrac{a}{c}\)

OpenStudy (jdoe0001):

the pythagorean theorem, doesn't tell us what the value should be, negative or minus, so \(\bf a = \pm \sqrt{3^2-2^4}\implies a = \pm \sqrt{5}\)

OpenStudy (jdoe0001):

negative or positive I meant

OpenStudy (anonymous):

Okay thank you I have the rest of the equation figured out

OpenStudy (jdoe0001):

yw

OpenStudy (jdoe0001):

hmm 2^4? hehe, anyhow, I meant \(\bf a = \pm \sqrt{3^2-2^2}\implies a = \pm \sqrt{5}\)